# KAG-Thinker: Teaching Large Language Models to Think with Human-like Reasoning Process

### v1

@@ 4.2 Main Results
We evaluated our framework on seven widely used benchmark datasets. Here, to maintain consistency with the retriever components of other baseline methods, we conduct plain text retrieval using the "Step" content from the logical form. Table 6 presents the comparisons between our model and other baselines. Compared to baselines without retrieval, the performance of our model exceeds Naive Generation and CoT by an average of 27. 1% and 34. 6%, respectively, in seven datasets. The primary reason for the improvement is that, within these datasets, retrieval is essential, as LLMs struggle to directly answer these questions accurately without access to relevant information. Compared to retrieval-augmented methods, our model outperforms Search-o1, IRCoT, and Naive RAG by average margins of 24.6%, 22.6%, and 14.8%, respectively. These enhancements are primarily attributed to our model’s implementation of breadth decomposition and depth solving, enabling it to learn how to leverage the retriever more effectively—particularly during deep retrieval, where the generated queries align well with the retriever’s capabilities. In comparison to reinforcement learning-based approaches, it can be observed that our model outperforms the previous state-of-the-art model, ReSearch, by an average of 4.1% in EM score across seven datasets. Specifically, it improves by an average of 4.5% on single-hop datasets and by an average of 3.9% on multi-hop datasets. The primary reason is that our breadth decomposition and depth solving framework enables the decomposition of questions into atomic granularity, thereby reducing the complexity involved in model retrieval and answering.
@@ 4.3 Thinker with KAG
To maintain consistency with KAG-V0.7 Liang et al. (2024), we continue to use its test set, with 1,000 examples each from HotpotQA, Musique, and 2Wiki (consistent with HippoRAG Gutierrez et al. (2024)). The data for our self-built corpus are entirely derived from the supporting facts of these three datasets, totaling 26,990 documents. Within the self-built corpus, and utilizing the same retriever and retrieved documents, our Thinker model still outperforms ReSearch and Search-R1 across three multi-hop datasets. For a fair comparison against the baselines, the original Thinker model could only perform pure text retrieval with BGE-M3 for natural language steps in its logical form, even for operations like Deduce and Math. However, the KAG framework robustly supports these operations. Table 7 illustrates that KAG-Thinker 7B, operating within the KAG framework, shows a marked improvement over Thinker, with average EM and F1 scores increasing by 3.0% and 3.8%, respectively. These enhancements are mainly attributable to KAG-Thinker 7B’s integration of KAG’s hybrid retriever (Sec 3.4) and KAG’s native support for Math and Deduce. This advancement allows KAG-Thinker 7B to achieve performance comparable to the larger KAG-V0.7 72B model, significantly cutting down on online inference demands and resource consumption.
| Framework | Retriever | HotpotQA | MusiQue | 2Wiki | Avg | ||||
@@
| KAG-V0.7 72B | KAG Hybrid Retriever | 0.574 | 0.706 | 0.326 | 0.458 | 0.657 | 0.741 | 0.519 | 0.635 |
| KAG-V0.7 32B | KAG Hybrid Retriever | 0.545 | 0.672 | 0.271 | 0.383 | 0.642 | 0.724 | 0.486 | 0.593 |
| KAG-V0.7 7B | KAG Hybrid Retriever | 0.443 | 0.557 | 0.181 | 0.276 | 0.419 | 0.501 | 0.348 | 0.445 |
| KAG-Thinker 7B | KAG Hybrid Retriever | 0.568 | 0.698 | 0.350 | 0.469 | 0.644 | 0.711 | 0.520 | 0.626 |
@@ 5 Medical Field Application
To address these issues, we design an interpretable framework. By decomposing problems into logically dependent sub-problems, the problem-solving logic becomes more rigorous and interpretable. Through knowledge boundary determination, it reduces the introduction of unnecessary noisy references, thereby improving the accuracy of the model’s answers.
@@ 5.1 Medical Thinker Model
We structure the samples into a multi-turn dialogue format. For synthesizing data, question decomposition is performed using Qwen2.5-72B-Instruct, which offers stronger reasoning capabilities and instruction-following abilities, allowing distillation into smaller models. The target model is trained to generate answers for sub-problems and determine whether external executors are required. Since the medical Q&A often involves lengthy answers for sub-problems, we call the LLM twice to confirm whether the answers are correct and complete, which serve as labels for the knowledge boundary determination. The subsequent deep retrieval, reasoning, and answer synthesis utilize distilled synthetic data derived from the LLMs.
5.2 Medical Knowledge Graph
@@ 5.2.1 Medical Data Sources
5.2.2 Medical Knowledge Graph Construction
For the extraction of information from the medical domain, we adopt schema constraint extraction due to the low error tolerance in the medical context. The use of a predefined schema restricts the extraction and generation to specific, structured key information (e.g., disease names, drug dosages, treatment plans), avoiding the introduction of irrelevant or erroneous content. We select 23 commonly used medical entity types that cover the main semantic categories involved in clinical diagnosis and treatment (e.g., diseases, symptoms, medications, diagnostic tests) and extend to general concepts related to the medical domain, public health, and system integration (e.g., disciplines, organizations, biological processes). This classification system refers to international medical terminology standards (for example, SNOMED CT, ICD-10) and ontologies of representative medical knowledge graphs, which effectively support subsequent structured knowledge fusion, reasoning, and generation tasks. Specific medical entity types can be found in Appendix 14. The detailed structured knowledge extraction process and prompt references can be found in Appendix 15. Figure 7 illustrates the constructed medical knowledge graph and shows the connectivity of the structured example information.

### v2

@@ 4.2 Main Results
We evaluated our model on seven widely used benchmark datasets. Here, to maintain consistency with the retriever components of other baseline methods, we conduct plain text retrieval using the "Step" content from the logical form. Table 6 presents the comparisons between our model and other baselines. Compared to baselines without retrieval, the performance of our model exceeds Naive Generation and CoT by an average of 27. 1% and 34. 6%, respectively, in seven datasets. The primary reason for the improvement is that, within these datasets, retrieval is essential, as LLMs struggle to directly answer these questions accurately without access to relevant information. Compared to retrieval-augmented methods, our model outperforms Search-o1, IRCoT, and Naive RAG by average margins of 24.6%, 22.6%, and 14.8%, respectively. These enhancements are primarily attributed to our model’s implementation of breadth decomposition and depth solving, enabling it to learn how to leverage the retriever more effectively—particularly during deep retrieval, where the generated queries align well with the retriever’s capabilities. In comparison to reinforcement learning-based approaches, it can be observed that our model outperforms the previous state-of-the-art model, ReSearch, by an average of 4.1% in EM score across seven datasets. Specifically, it improves by an average of 4.5% on single-hop datasets and by an average of 3.9% on multi-hop datasets. The primary reason is that our breadth decomposition and depth solving model enables the decomposition of questions into atomic granularity, thereby reducing the complexity involved in model retrieval and answering.
@@ 4.3 Thinker with KAG
We create KAG-Thinker by applying the Thinker model within the KAG framework. The multi-turn interactive reasoning framework of KAG-Thinker is depicted in Figure 6. To maintain consistency with KAG Liang et al. (2024), we continue to use its test set, with 1,000 examples each from HotpotQA, Musique, and 2Wiki (consistent with HippoRAG Gutierrez et al. (2024)). The data for our self-built corpus are entirely derived from the supporting facts of these three datasets, totaling 26,990 documents. Within the self-built corpus, and utilizing the same retriever and retrieved documents, our Thinker model still outperforms ReSearch and Search-R1 across three multi-hop datasets.
| Methods | Retriever | HotpotQA | MusiQue | 2Wiki | Avg | ||||
@@
| KAG-V0.7 72B | Hybrid Graph Retriever | 0.574 | 0.706 | 0.326 | 0.458 | 0.657 | 0.741 | 0.519 | 0.635 |
| KAG-V0.7 32B | Hybrid Graph Retriever | 0.545 | 0.672 | 0.271 | 0.383 | 0.642 | 0.724 | 0.486 | 0.593 |
| KAG-V0.7 7B | Hybrid Graph Retriever | 0.443 | 0.557 | 0.181 | 0.276 | 0.419 | 0.501 | 0.348 | 0.445 |
| KAG-Thinker 7B | Hybrid Graph Retriever | 0.568 | 0.698 | 0.350 | 0.469 | 0.644 | 0.711 | 0.520 | 0.626 |
For a fair comparison against the baselines, the original Thinker model could only perform pure text retrieval with BGE-M3 for natural language steps in its logical form, even for operations like Deduce and Math. However, the KAG framework robustly supports these operations. Table 7 illustrates that KAG-Thinker 7B, operating within the KAG framework, shows a marked improvement over Thinker, with average EM and F1 scores increasing by 3.0% and 3.8%, respectively. These enhancements are mainly attributable to KAG-Thinker 7B’s integration of KAG’s hybrid graph retriever (Sec 3.4) and KAG’s native support for Math and Deduce. This advancement allows KAG-Thinker 7B to achieve performance comparable to the larger KAG-V0.7 72B model, significantly cutting down on online inference demands and resource consumption.
@@ 5 Medical Field Application
To address these issues, we design an interpretable model. By decomposing problems into logically dependent sub-problems, the problem-solving logic becomes more rigorous and interpretable. Through knowledge boundary determination, it reduces the introduction of unnecessary noisy references, thereby improving the accuracy of the model’s answers.
@@ 5.1 Medical Thinker Model
We structure the samples into a multi-turn interactive reasoning format. For synthesizing data, question decomposition is performed using Qwen2.5-72B-Instruct, which offers stronger reasoning capabilities and instruction-following abilities, allowing distillation into smaller models. The target model is trained to generate answers for sub-problems and determine whether external executors are required. Since the medical Q&A often involves lengthy answers for sub-problems, we call the LLM twice to confirm whether the answers are correct and complete, which serve as labels for the knowledge boundary determination. The subsequent deep retrieval, reasoning, and answer synthesis utilize distilled synthetic data derived from the LLMs.
5.2 Medical Knowledge Base
@@ 5.2.1 Medical Data Sources
5.2.2 Medical Knowledge Base Construction
For the extraction of information from the medical domain, we adopt schema constraint extraction due to the low error tolerance in the medical context. The use of a predefined schema restricts the extraction and generation to specific, structured key information (e.g., disease names, drug dosages, treatment plans), avoiding the introduction of irrelevant or erroneous content. We select 23 commonly used medical entity types that cover the main semantic categories involved in clinical diagnosis and treatment (e.g., diseases, symptoms, medications, diagnostic tests) and extend to general concepts related to the medical domain, public health, and system integration (e.g., disciplines, organizations, biological processes). This classification system refers to international medical terminology standards (for example, SNOMED CT, ICD-10) and ontologies of representative medical knowledges, which effectively support subsequent structured knowledge fusion, reasoning, and generation tasks. Specific medical entity types can be found in Appendix 14. The detailed structured knowledge extraction process and prompt references can be found in Appendix 15. Figure 8 illustrates the constructed medical knowledge base and shows the connectivity of the structured example information.
@@ 7 Contributors
Special acknowledgments.
Haofen Wang@Tongji University
Huajun Chen@Zhejiang University
Zhejiang University-Ant Group KG Joint Laboratory